OpenStudy (anonymous):

what is the graph of the equation (x+8)^2+(y-1)^2=9

OpenStudy (jdoe0001):

notice that \(\bf 9 = 3^2\) so \(\bf (x+8)^2+(y-1)^2=9 \implies (x+8)^2+(y-1)^2=3^2\) http://www.mathwarehouse.com/geometry/circle/images/equation-of-circle/general-formula-equation-of-circle.png